TETSXPREME:

A net force of 60 N north acts on an object with a mass of 30 kg. Use Newton's second law of motion to calculate the amount of acceleration the object will experience. Then explain how the amount of acceleration will change if the net force or the mass of the object increases. @vocaloid

TETSXPREME:

60=30a 60/30=30 a/30 2=1a a=2m/s^2

Joe348:

F=M X A

Joe348:

60=30 times mass

TETSXPREME:

If the mass increases then the acceleration decreases. If force increase then acceleration increases

Joe348:

60/30=2

TETSXPREME:

I just wanted to know if it was right or not :)

Joe348:

acceleration is 2 m/s^2

YRJ8498:

F = m×a60 N = 30 kg × a30 × a = 60a = 60/30a = 2 m/s²

YRJ8498:

Acceleration depends directly on the net force acting on an object, and inversely on its mass. The amount of acceleration increases if the net force increases while the object's mass remains the same. When the object's mass increases while the net force remains the same, the amount of acceleration decreases.